Calculate 🔺E if 100kJ of heat energy is absorbed by the system and 30kJ of work is done on the surroundings.


1
Expert's answer
2021-05-13T02:08:29-0400

As given that work is done on the surroundings and heat is absorbed by the system.

Therefore,

q = 100kJ

w = −30kJ

Now, from first law of thermodynamics,

Change in internal energy-


ΔU = q + w

⇒ ΔU = 100 + (−30) = 70kJ



Hence, ΔE is 70kJ